You decided to put away 5%

 of your annual salary for

 retirement. Your annual

 salary is $20,000 for the first

 year (at age 22) and it

 increases by 3% every year

 until your retirement (at age

 50). What will be the worth of

 your account when you retire

 if the interest is 4.5% per

 year?


1
Expert's answer
2021-08-17T16:56:34-0400

We use the formula below:

"A=P(1+\frac {r}{n})^{nt}"

"P= \frac {5}{100}\times 20,000=1,000"

"r=4.5" %

"t=50-22=28"

"n=0.03"

"A= 1,000(1+\frac {4.5}{0.03})^{2.52}"

"A=309,757,554.67"